Grade 12Mechanics

an aeroplane flying horizontally 1km above the ground is observed at an elevation of 60degree and after 10s the elevation is observed to be 30degree. The uniform speed of the aeroplane in kilometer per hr is

s
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
Tan60°= 1/d =√3
⇒d =1/√3 km 
now, the aeroplane cover (d+x) km
tan30°= 1/(d+x)=1/√3
⇒(d+x) = √3
⇒x = √3 -(1/√3) = (3-1)/√3 =2/√3= (2×1.732)/3 = 1.15 km
so speed =distance/time = 1.15/(10/3600) =414 km/hr